OpenStudy (anonymous):

Need answers checked What is y = x2 – 12x + 20 written in vertex form? y =(x – 12)2 + 36 y =(x – 12)2 + 26 y =(x – 6)2 – 16 y =(x – 6)2 + 20 An open box is formed by cutting squares with side lengths of 3 inches from each corner of a square piece of paper. What is a side length of the original paper if the box has a volume of 675 cubic inches? 21 inches 24 inches 33 inches 36 inches

OpenStudy (anonymous):

D for the first one A for the second

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

The first one is C y=x2-12x+20 y-20=x2-12x, with that you wanna make it a perfect square on the right side. 36 will do that. y-20+36=x2-12x+36 y+16=(x-6)^2 y=(x-6)^2 -16

OpenStudy (anonymous):

A is correct for the second one. Bruhhh is correct for the first one.